OpenStudy (asad786):

WILL MEDAL Which expression is equivalent to this one? –7(5 + 10) A. (–7) · 5 + (–7) · 10 B. (–7) · 5 + 10 C. (–7) + 5 + 10 D. (–7) · 5 · 10

OpenStudy (asad786):

i think B

OpenStudy (igreen):

No, when we distribute we multiply the number outside of the parenthesis to every number inside the parenthesis..

OpenStudy (igreen):

So what do you think?

OpenStudy (swamijatin2):

i think it D

OpenStudy (asad786):

same

OpenStudy (igreen):

@swamijatin2 I just told you not to give answers -.-

OpenStudy (igreen):

And that's incorrect.

OpenStudy (igreen):

What numbers are in the parenthesis? @Asad786

OpenStudy (asad786):

1 sec

OpenStudy (swamijatin2):

i m just geusing

OpenStudy (asad786):

-7

OpenStudy (swamijatin2):

i know

OpenStudy (igreen):

Inside, not outside

OpenStudy (swamijatin2):

can i tell

OpenStudy (igreen):

No

OpenStudy (asad786):

5 + 10

OpenStudy (igreen):

Yes, 5 and 10..and -7 is outside..so which one multiplies -7 to 5 and 10?

OpenStudy (swamijatin2):

x(1+2) 1x(X) +1x2

OpenStudy (igreen):

\(\sf a(b + c) = ab + ac\)

OpenStudy (asad786):

a

OpenStudy (igreen):

Correct

OpenStudy (asad786):

thanks

OpenStudy (swamijatin2):

ya same

OpenStudy (igreen):

\(\sf -7(5 + 10) \rightarrow -7 \cdot 5 + -7 \cdot 10\)